# ICP-4D: Bridging Iterative Closest Point and LiDAR Panoptic Segmentation

### v1

@@ 5 Conclusion
In this paper, we introduce a novel instance association strategy ICP-4D for 4D LiDAR panoptic segmentation, the first training-free framework tailored for a single scan setting. Our proposed Sinkhorn-based soft matching forming instance-aware correspondence enables ICP procedure to precisely align between two consecutive instance point segments, thereby improving temporal continuity. Additionally, we systematically design overall pipeline, underscoring the instance types to alleviate the inherent challenges of ICP and LiDAR sensors. Through experimental evaluations on both SemanticKITTI and nuScenes, our ICP-4D showcases its superiority not only on the validation sets but also on the hidden test sets, demonstrating robustness across various sensor configuration. Ultimately, our approach offers a promising direction for 4D perception by solely leveraging strong 3D panoptic models to support temporal understanding.
Limitations. Although ICP-4D shows substantial promise in advancing 4D perception, several limitations persist that warrant further investigation, particularly in real-time deployment, handling partial LiDAR observations, and dependence on camera parameters. We provide additional discussions in the supplementary materials.
Supplementary Material
@@ A.1 Baseline Details
In this section, we provide a detailed overview of the strong baselines considered in our comparison: 4D-PLS [2] (CVPR’21), 4D-Stop (ECCVW’22) [20], EQ-4D-Stop (ICCV’23) [63], CA-Net (R-AL’22) [29], Mask4D (R-AL’23) [30], and Mask4Former (ICRA’24) [54]. Note that we report the official results as stated in the original papers, and reproduce the results when formal reports are not available. The following paragraph details the experimental settings and notable modifications for each baseline.
@@ A.2 Metrics
In this section, we provide detailed explanations of the metrics employed for evaluating association performance in our experiments. While LSTQ is the most widely used metric, we additionally report two complementary metrics, MOTSA and PTQ, to provide a more comprehensive assessment.
LSTQ [2] is delicately designed to evaluate 4D panoptic segmentation quality. First, to assess the semantic segmentation performance over the classes $C$, the classification score Scls is defined as:
@@
MOTSA & sMOTSA [45] are mask-based variants of the widely used Multi-Object Tracking Accuracy (MOTA) metric, replacing bounding-box overlaps with mask IoU to better evaluate segmentation-level tracking performance. Formally,
@@
PTQ & sPTQ [17] are formulated to satisfy the fundamental requirement that panoptic tracking metrics must jointly capture segmentation accuracy and tracking consistency over time. Thus, they incorporate an ID-switch penalty into the original Panoptic Quality (PQ) measure, as follows:
@@
Algorithm 1 Instance association process of ICP-4D
0:  Instance predictions $\mathcal{I}$, semantic predictions $\mathcal{S}$, memory bank $\mathcal{B}$
1:  for time step $t\in\{1,2,\dots,T-1\}$ do
2:   Construct the candidate set $\mathcal{M}=\mathcal{G}^{\prime}_{\text{src}}\times\mathcal{G}^{\prime}_{\text{dst}}$ filtered by semantic predictions $\mathcal{S}$. // Static instances (Sec. 3.2.1)
3:   Associate instances using the center $\mu$ of each point set (Eq. 4).
4:   Refine the matched pairs using the covariance descriptor $\sigma$ (Eq. 6). // Dynamic instances (Sec. 3.2.2)
5:   Perform dynamic instance association using ICP enhanced with Sinkhorn-based soft matching. // Missing instances (Sec. 3.2.3)
@@
7:    Apply the dynamic matching procedure (Lines 5) to missing instances.
@@
10:  Consistent instance IDs across the entire sequence.
@@ B.1 Pipeline Details
The overall association pipeline of our proposed ICP-4D is summarized in Algorithm 1. We apply ICP-4D sequentially across all scans within each sequence, and reset the memory bank $\mathcal{B}$ at the beginning of every new sequence.
@@ B.2 Complexity Analysis
In this section, we provide a computational complexity comparison between Sinkhorn-based soft matching and nearest neighbor-based matching. As described in Section.3.2.2, we utilize the entropy-regularized optimal transport problem [12] to obtain instance-aware point correspondences. Before analyzing the regularized form, we briefly review the optimal transport problem. Given two probability vectors $r\in\mathbb{R}_{+}^{n}\;\text{and}\;c\in\mathbb{R}_{+}^{m}$, the original optimal transport problem [44] aims to find the optimal transport plan $\mathcal{Q}\in U(r,\;c)$. The entries of $\mathcal{Q}$ represent the amount of mass transported from the $i$-th element of $r$ to the $j$-th element of $c$. The transport polytope $U$ is defined as:
$U(r,\;c)$ $=$ (15)
$\{Q\in$ $\mathbb{R}^{n\times m}_{+}|\sum_{j}Q_{ij}=r_{i},\;\forall i,\;\sum_{i}Q_{ij}=c_{j},\;\forall j\}.$
@@
This optimization is a linear program, for which the network simplex [3] is the most widely used exact solver. In the symmetric case $n=m$, its worst-case computational complexity scales as $O(n^{3}\log n)$. However, due to its heavy computational cost, adding an entropic regularization is widely adopted to accelerate optimal transport. By incorporating the Shannon entropy $H(\mathcal{Q}):=-\sum_{i,j}\mathcal{Q}_{ij}\log\mathcal{Q}_{ij}$ into Eq. 16, we obtain the entropy-regularized optimal transport formulation. To handle marginal constraints, we introduce dual variables $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ and form the corresponding Lagrangian. Then, by taking the first-order optimality condition, we find that optimal solution $\mathcal{Q}$ has matrix form $Q=\text{diag}(u)K\text{diag}(v)$, where $u_{i}=e^{f_{i}/\epsilon}$, $v_{j}=e^{g_{j}/\epsilon}$, and $K_{ij}=e^{-Z_{ij}/\epsilon}$. Since $u$ and $v$ must satisfy the marginal constraints, the entropic regularization transforms the classical optimal transport problem from a linear program into a matrix scaling problem. Thus, when $n=m$, the computational complexity is reduced to $O(n^{2})$. However, the nearest neighbor algorithm operates in $O(n\log n)$ on average and up to $O(n^{2})$ in the worst case in practice, as implemented in the scikit-learn library (e.g., using KD-Tree [5]). Consequently, our Sinkhorn-based solution is theoretically comparable to the nearest neighbor algorithm, and the wall-clock time shows only a marginal difference while achieving higher performance (see Table A.1).
Table A.1: Comparison Between Nearest Neighbor and Sinkhorn-based Matching. We report the results of our full ICP-4D when integrating each correspondence strategy.
w/ Nearest Neighbor (KD-Tree) w/ Sink.-based Soft Matching
LSTQ Inf. Time (s) LSTQ Inf. Time (s)
72.1 2.01 72.7 2.35
Table A.2: SemanticKITTI [4] Test Set Leaderboard (Top 5). mAQ and mIoU denote the Sassoc and Scls, respectively.
@@ Appendix C Additional Experimental Results
black!10black 4D-Stop [20] 2 15.9 3.67 66.4 71.8 61.4 64.9 64.1 62.9 67.3 58.8 68.3 53.3
@@
black!10black Eq-4D-Stop [63] 2 11.3 2.89 68.9 74.8 63.5 65.7 68.4 65.4 69.7 61.5 68.7 59.1
@@
black!10black Mask4Former [54] 2 10.4 3.51 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
@@ C.1 Leaderboard Results
Table A.2 shows the SemanticKITTI [4] challenge leaderboard ranks on the hidden test set, ordered by LSTQ. In this leaderboard, we observe that our proposed ICP-4D outperforms not only all published methods but also every competition participant. Specifically, we achieve 1st place in LSTQ as of submission date. These results demonstrate the strong generalization capability and practical effectiveness of our method under challenging scenarios.
C.2 Additional Results
Table A.4: Additional comparisons across multiple metrics. ( ) denotes the number of input scans (i.e., 1f = single, 2f = two).
Method Metric MOTSA sMOTSA PTQ sPTQ
4D-StOP (2f) [20] 31.0 27.1 53.4 53.7
EQ-4D-StOP (2f) [63] 43.6 39.6 56.4 56.7
Mask4D (1f) [30] 52.1 48.1 58.4 58.5
Mask4Former (2f) [54] 47.3 42.1 48.3 48.5
black!10black Mask4Former (1f) - - - -
@@ C.2 Additional Results
[…]
Table A.5: Comparison of the quantitative results on panoptic nuScenes [13].
@@
Mask4Former [54] 67.6 63.0 72.5 48.3/76.9 42.6/37.0 60.5/90.2 76.7/92.7 42.7/44.9 61.1/77.0 53.5/74.0 48.9/55.7 38.7/47.8 62.2/77.8 95.6 69.2 71.5 74.6 88.5 87.0
Mask4D [30] 58.5 54.1 63.1 29.7/64.3 27.8/20.5 60.5/79.2 72.5/83.5 29.5/21.8 52.5/72.0 42.3/65.9 39.5/43.0 19.2/35.2 58.9/71.8 94.2 53.9 64.9 71.3 85.2 84.5
CA-Net [29] 63.4 56.6 71.0 30.1/74.5 31.9/29.1 56.5/88.5 73.3/91.2 38.6/43.2 63.2/78.7 53.6/71.2 34.6/54.6 39.5/48.3 62.1/77.3 95.6 68.4 70.3 72.7 86.7 85.2
ICP-4D (Ours) 67.9 65.0 71.0 40.7/74.5 44.1/29.1 58.9/88.5 80.9/91.2 40.1/43.2 65.9/78.7 67.2/71.2 52.0/54.6 40.6/48.3 65.7/77.3 95.6 68.4 70.3 72.7 86.7 85.2
Results on Multiple Metrics. We further present additional results with multiple metrics to enable a more comprehensive comparison. MOTSA [45] is widely used for evaluating multi-object tracking and segmentation. Although MOTSA emphasizes the recognition quality rather than temporal consistency, it nonetheless offers a straightforward indicator of the ID switches. Moreover, PTQ [17] is designed to jointly evaluate panoptic segmentation and tracking quality in a holistic manner. In Table C.2, evaluations are carried out on the SemanticKITTI [4] validation set. We observe that ICP-4D consistently surpasses all baselines across multiple metrics, including LSTQ. This indicates that the strength of ICP-4D goes beyond any particular metric, offering a robust and practically reliable solution for instance association.
C.3 Hyperparameter Analysis
In this section, we examine the impact of different matching thresholds on instance association performance. Figure A.1 illustrates the LSTQ scores of ICP-4D on the SemanticKITTI [4] validation set. We empirically determine the matching thresholds $\tau_{iou}$, $\tau_{dist}$, $\tau_{center}$, $\tau_{cov}$} over {0.1, 0.2, 0.3, 0.4, 0.5}. Here, we observe three insights. First, extremely low value of threshold $\tau_{iou}$ often yields decent performance due to over-association, whereas high thresholds lead to over-filtering and degraded matching quality. Second, excessively large $\tau_{center}$ and $\tau_{cov}$ cause static instances to be matched too liberally, leading to unreliable association. Lastly, our proposed ICP-4D delivers performance gains over all baselines, regardless of the hyperparameter choices.
Figure A.1: Hyperparameter Analysis.
C.4 Quantitative Results

### v2

@@ 5 Conclusion
In this paper, we introduce a novel instance association strategy Geo-4D for 4D LiDAR panoptic segmentation, the first training-free framework tailored for a single scan setting. Our proposed Global Geometry-aware Soft Matching forming instance-aware correspondence enables ICP procedure to precisely align between two consecutive instance point segments, thereby improving temporal continuity. Additionally, we systematically design the overall pipeline, underscoring the instance types to alleviate the inherent challenges of ICP and LiDAR sensors. Through experimental evaluations on both SemanticKITTI and nuScenes, our Geo-4D showcases its superiority not only on the validation sets but also on the hidden test sets, demonstrating robustness across various sensor configurations. Ultimately, our approach offers a promising direction for 4D perception by solely leveraging strong 3D panoptic models to support temporal understanding.
Limitations. Although Geo-4D shows substantial promise in advancing 4D perception, several limitations persist that warrant further investigation, particularly in real-time deployment, handling partial LiDAR observations, and dependence on accurate ego pose estimation. We provide additional discussions in the supplementary materials.
@@ A.1 Baseline Details
In this section, we provide a detailed overview of the strong baselines considered in our comparison: 4D-PLS [Aygun_2021_CVPR] (CVPR’21), 4D-Stop (ECCVW’22) [kreuzberg20224d], EQ-4D-Stop (ICCV’23) [zhu20234d], CA-Net (R-AL’22) [marcuzzi2022contrastive], Mask4D (R-AL’23) [marcuzzi2023mask4d], and Mask4Former (ICRA’24) [yilmaz2024mask4former]. Note that we report the official results as stated in the original papers, and reproduce the results when formal reports are not available. The following paragraph details the experimental settings and notable modifications for each baseline.
@@ A.2 Metrics
We provide detailed explanations of the metrics employed for evaluating association performance in our experiments. While LSTQ is the most widely used metric, we additionally report two complementary metrics, MOTSA and PTQ, to provide a more comprehensive assessment.
LSTQ [Aygun_2021_CVPR] is delicately designed to evaluate 4D panoptic segmentation quality. First, to assess the semantic segmentation performance over the classes $C$, the classification score Scls is defined as:
@@
MOTSA & sMOTSA [voigtlaender2019mots] are mask-based variants of the widely used Multi-Object Tracking Accuracy (MOTA) metric, replacing bounding-box overlaps with mask IoU to better evaluate segmentation-level tracking performance. Formally,
@@
PTQ & sPTQ [hurtado2020mopt] are formulated to satisfy the fundamental requirement that panoptic tracking metrics must jointly capture segmentation accuracy and tracking consistency over time. Thus, they incorporate an ID-switch penalty into the original Panoptic Quality (PQ) measure, as follows:
@@
Algorithm 1 Overall pipeline of Geo-4D
0: Panoptic prediction $(s,l)$, memory bank $\mathcal{B}$
1:for sequence $t\in\{1,2,\dots,T-1\}$ do
2:  Construct the candidate set $\mathcal{M^{\prime}}$ filtered by semantic predictions $s$. // Static instances
3:  Associate instances using the centroid $\mu$ of each point set (Eq. 5).
4:  Refine the matched pairs using the covariance descriptor $\sigma$ (Eq. 7). // Dynamic instances
5:  Perform global geometric association using ICP enhanced with global geometry-aware soft matching. // Missing instances
@@
7:   Apply the dynamic matching procedure (Line 5) to missing instances.
@@
10: Consistent instance IDs throughout the entire sequence.
@@ B.1 Pipeline Details
The overall association pipeline of our proposed Geo-4D is summarized in Algorithm 1. We apply Geo-4D sequentially across all scans within each sequence, and reset the memory bank $\mathcal{B}$ at the beginning of every new sequence.
Table A.1: Comparison Between Nearest Neighbor and Global Geometry-aware Soft Matching (GGSM). We report the results of our full Geo-4D when integrating each correspondence strategy.
w/ Nearest Neighbor w/ GGSM (Ours)
LSTQ Inf. Time (s) LSTQ Inf. Time (s)
72.1 2.01 72.7 2.35
@@ B.2 Complexity Analysis
In this section, we provide a computational complexity comparison between Global Geometry-aware Soft Matching (GGSM) and nearest neighbor-based matching. As described in the main manuscript, we utilize the entropy-regularized optimal transport problem [cuturi2013sinkhorn] to obtain instance-aware point correspondences. Before analyzing the regularized form, we briefly review the optimal transport problem. Given two probability vectors $r\in\mathbb{R}_{+}^{n}\;\text{and}\;c\in\mathbb{R}_{+}^{m}$, the original optimal transport problem [villani2008optimal] aims to find the optimal transport plan $\mathcal{Q}\in U(r,\;c)$. The entries of $\mathcal{Q}$ represent the amount of mass transported from the $i$-th element of $r$ to the $j$-th element of $c$. The transport polytope $U$ is defined as:
$U(r,\;c)=\{Q\in\mathbb{R}^{n\times m}_{+}|\sum_{j}Q_{ij}=r_{i},\;\forall i,\;\sum_{i}Q_{ij}=c_{j},\;\forall j\}.$ (16)
@@
This optimization is a linear program, for which the network simplex [bazaraa2011linear] is the most widely used exact solver. In the symmetric case $n=m$, its worst-case computational complexity scales as $O(n^{3}\log n)$. However, due to its heavy computational cost, adding an entropic regularization is widely adopted to accelerate optimal transport. By incorporating the Shannon entropy $H(\mathcal{Q}):=-\sum_{i,j}\mathcal{Q}_{ij}\log\mathcal{Q}_{ij}$ into Eq. 17, we obtain the entropy-regularized optimal transport formulation. To handle marginal constraints, we introduce dual variables $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ and form the corresponding Lagrangian. Then, by taking the first-order optimality condition, we find that optimal solution $\hat{\mathcal{Q}}$ has matrix form $\hat{\mathcal{Q}}=\text{diag}(u)K\text{diag}(v)$, where $u_{i}=e^{f_{i}/\epsilon}$, $v_{j}=e^{g_{j}/\epsilon}$, and $K_{ij}=e^{-Z_{ij}/\epsilon}$. Since $u$ and $v$ must satisfy the marginal constraints, the entropic regularization transforms the classical optimal transport problem from a linear program into a matrix scaling problem. Thus, when $n=m$, the computational complexity is reduced to $O(n^{2})$. However, the nearest neighbor algorithm operates in $O(n\log n)$ on average and up to $O(n^{2})$ in the worst case in practice, as implemented in the scikit-learn library (e.g., using KD-Tree [bentley1975multidimensional]). Consequently, our GGSM solution is theoretically comparable to the nearest neighbor algorithm, and the wall-clock time shows only a marginal difference while achieving higher performance (see Table A.1).
B.3 Option: Bijective assignment.
As mentioned in the main manuscript, our simple heuristic matching strategy consistently outperforms more complex assignment schemes, despite its simplicity. Here, we provide additional comparisons and describe an alternative bijective assignment formulation. A naive approach in global geometric association maintains consistent instance identities by assigning a single instance ID to matched candidates; however, this fails to enforce one-to-one correspondence. Thus, we derive a transformation-aware matching cost that integrates the global geometric association results—rotation $\mathbf{R}$, translation $\mathbf{t}$, and the IoU score $s_{\text{iou}}$. Formally, the matching cost $\mathcal{C}$ of dynamic matching pairs $(i,j)\in\mathcal{M}^{\prime\prime}$ is defined as:
$c_{ij}^{(t)}=\|\mathbf{t}_{ij}\|_{2},$ $c_{ij}^{(r)}=\|\theta_{ij}\|_{2},\quad c_{ij}^{(s)}=\|1-s_{iou}\|_{2},$
where $\theta$ denotes the axis–angle rotation vector obtained from $\mathbf{R}$. We normalize each cost term to comparable scales before weighting so that $\gamma_{t}$, $\gamma_{r}$, and $\gamma_{s}$ control their relative contributions. By minimizing $\mathcal{C}$ with the Hungarian algorithm [kuhn1955hungarian], we obtain a bijective assignment that considers the geometric alignment. We present experimental results in Section C.3.
Table A.2: SemanticKITTI [behley2019semantickitti] Test Set Leaderboard (Top 5). mAQ and mIoU denote Sassoc and Scls, respectively.
@@ Appendix C Additional Experimental Results
black!10black 4D-Stop [kreuzberg20224d] 2 15.9 3.67 66.4 71.8 61.4 64.9 64.1 62.9 67.3 58.8 68.3 53.3
@@
black!10black Eq-4D-Stop [zhu20234d] 2 11.3 2.89 68.9 74.8 63.5 65.7 68.4 65.4 69.7 61.5 68.7 59.1
@@
black!10black Mask4Former [yilmaz2024mask4former] 2 10.4 3.51 70.5 74.3 66.9 67.1 66.6 68.4 67.3 69.6 72.7 65.3
@@ C.1 Leaderboard Results
Table A.2 shows the SemanticKITTI [behley2019semantickitti] challenge leaderboard ranks on the hidden test set, ordered by LSTQ. In this leaderboard, we observe that our proposed Geo-4D outperforms not only all published methods but also every competition participant. Specifically, we achieve 1st place in LSTQ as of submission date. These results demonstrate the strong generalization capability and practical effectiveness of our method under challenging scenarios.
C.2 Effect of the Number of Scans.
@@
Table A.4: Comparison of the quantitative results on panoptic nuScenes [fong2022panoptic].
@@
Mask4Former [yilmaz2024mask4former] 67.6 63.0 72.5 48.3/76.9 42.6/37.0 60.5/90.2 76.7/92.7 42.7/44.9 61.1/77.0 53.5/74.0 48.9/55.7 38.7/47.8 62.2/77.8 95.6 69.2 71.5 74.6 88.5 87.0
Mask4D [marcuzzi2023mask4d] 58.5 54.1 63.1 29.7/64.3 27.8/20.5 60.5/79.2 72.5/83.5 29.5/21.8 52.5/72.0 42.3/65.9 39.5/43.0 19.2/35.2 58.9/71.8 94.2 53.9 64.9 71.3 85.2 84.5
CA-Net [marcuzzi2022contrastive] 63.4 56.6 71.0 30.1/74.5 31.9/29.1 56.5/88.5 73.3/91.2 38.6/43.2 63.2/78.7 53.6/71.2 34.6/54.6 39.5/48.3 62.1/77.3 95.6 68.4 70.3 72.7 86.7 85.2
Geo-4D (Ours) 67.9 65.0 71.0 40.7/74.5 44.1/29.1 58.9/88.5 80.9/91.2 40.1/43.2 65.9/78.7 67.2/71.2 52.0/54.6 40.6/48.3 65.7/77.3 95.6 68.4 70.3 72.7 86.7 85.2
Table A.5: Ablation on matching strategy.
Method LSTQ $_{\text{assoc}} S$$_{\text{cls}}$
Hungarian Matching 72.3 77.6 67.5
Greedy Matching 72.7 78.3 67.5
C.3 Analysis of Matching Strategy.
[…]
C.4 Analysis of Global Geometry-aware Soft Matching.
In the main paper, we argue that the proposed Global Geometry-aware Soft Matching (GGSM) improves the robustness of the matching strategy against noisy measurements and predictions. To isolate the contribution of GGSM, we additionally evaluate its effectiveness on non-rigid point cloud registration benchmarks [li2022lepard, li2022non] independent of tracking performance. This allows us to analyze whether GGSM improves geometric correspondence estimation under challenging conditions such as partial overlap, occlusion, and large motion.
Baselines. While more advanced baselines exist for non-rigid point cloud registration, we compare against several representative registration methods to better understand the effectiveness of GGSM. Specifically, we include Iterative Closest Point (ICP) [besl1992method] as a classical baseline, Sinkhorn-based registration which is closely related to our global geometry association, Non-rigid ICP (NICP) [newcombe2015dynamicfusion] as an extension of ICP for non-rigid registration, and Neural Scene Flow Prior (NSFP) [li2021neural] which predicts point-wise motion using neural networks. ICP determines the global rotation and translation parameters, which are designed for rigid transformations. Optimal Transport (OT) [cuturi2013sinkhorn] iteratively estimates point displacements by optimizing a transport plan via the Sinkhorn-Knopp algorithm [sinkhorn1967diagonal] and updating point locations according to the gradients of the OT objective. NICP estimates the movement of all points individually rather than relying on a single rigid transformation. We adopt a typical solution that optimizes a deformation graph with regularization constraints to enforce smooth deformations. NSFP optimizes a multilayer perceptron (MLP) to predict point-wise motion by minimizing the Chamfer distance between two point clouds.
Table A.6: Quantitative comparison on 4DMatch-F [li2022non] and 4DLoMatch-F [li2022non].
4DMatch-F 4DLoMatch-F
[…]
ICP [besl1992method] 0.296 2.96 12.06 71.50 0.565 0.14 0.74 90.87 0.10
Sinkhorn [sinkhorn1967diagonal] 0.308 2.76 8.13 79.86 0.505 0.20 0.86 89.47 3.76
NICP [newcombe2015dynamicfusion] 0.325 6.44 12.10 80.40 0.517 0.18 0.73 92.37 4.80
NSFP [li2021neural] 0.265 8.66 18.65 64.96 0.495 0.38 1.56 84.77 39.54
Ours (ICP+GGSM) 0.210 2.51 11.26 62.58 0.431 0.18 1.19 78.80 1.62
Runtime measured on a single NVIDIA A100.
[…]
Experimental Analysis. As shown in Tab. A.6, two key observations emerge across both 4DMatch-F and 4DLoMatch-F. First, GGSM achieves the lowest EPE and Outlier rates despite relying only on rigid transformations. Although non-rigid methods achieve higher point-wise accuracy (AccS and AccR), they often introduce globally inconsistent deformations, which can degrade the overall geometric alignment. In contrast, GGSM leverages global geometric structure to estimate a coherent transformation, leading to more reliable overall registration performance. Second, GGSM provides a well-balanced trade-off between efficiency and performance. NSFP models point-wise motion by optimizing a neural network for each point coordinate, which offers strong flexibility but incurs significant computational overhead. Moreover, NICP estimates per-vertex deformation parameters through iterative non-rigid optimization, which also limits efficiency. On the other hand, our GGSM remains computationally efficient as it inherits the efficiency of rigid ICP, resulting in strong overall registration quality by improving geometric consistency.
Takeaway: In this section, we validate the effect of GGSM in a controlled setting by excluding other factors (e.g., rigid objects and driving surface in on-road environments). It is important to note that the primary scope of this work is 4D LiDAR panoptic segmentation in driving scenes, rather than non-rigid point cloud registration itself. These experimental results support our claim that GGSM plays a crucial role in enabling robust and reliable instance association in driving scenes with both rigid and non-rigid instances. For additional discussions, please refer to the corresponding analyses presented in the main manuscript.
C.5 Quantitative Results